# DM-Codec: Distilling Multimodal Representations for Speech Tokenization

### v1

@@ Abstract
Recent advancements in speech-language models have yielded significant improvements in speech tokenization and synthesis. However, effectively mapping the complex, multidimensional attributes of speech into discrete tokens remains challenging. This process demands acoustic, semantic, and contextual information for precise speech representations. Existing speech representations generally fall into two categories: acoustic tokens from audio codecs and semantic tokens from speech self-supervised learning models. Although recent efforts have unified acoustic and semantic tokens for improved performance, they overlook the crucial role of contextual representation in comprehensive speech modeling. Our empirical investigations reveal that the absence of contextual representations results in elevated Word Error Rate (WER) and Word Information Lost (WIL) scores in speech transcriptions. To address these limitations, we propose two novel distillation approaches: (1) a language model (LM)-guided distillation method that incorporates contextual information, and (2) a combined LM and self-supervised speech model (SM)-guided distillation technique that effectively distills multimodal representations (acoustic, semantic, and contextual) into a comprehensive speech tokenizer, termed DM-Codec. The DM-Codec architecture adopts a streamlined encoder-decoder framework with a Residual Vector Quantizer (RVQ) and incorporates the LM and SM during the training process. Experiments show DM-Codec significantly outperforms state-of-the-art speech tokenization models, reducing WER by up to 13.46%, WIL by 9.82%, and improving speech quality by 5.84% and intelligibility by 1.85% on the LibriSpeech benchmark dataset. The code, samples, and model checkpoints are available at https://github.com/mubtasimahasan/DM-Codec.
@@ 1 Introduction
In recent years, the advent of Large Language Models (LLMs) has revolutionized various domains, offering unprecedented advancements across a wide array of tasks (OpenAI, 2024). A critical component of this success has been the tokenization of input data, enabling vast amounts of information processing (Du et al., 2024; Rust et al., 2021). Inspired by these breakthroughs, significant attention has shifted towards replicating similar successes in the realm of speech understanding and generation (Défossez et al., 2022; Hsu et al., 2021). However, tokenizing speech into discrete units presents unique challenges compared to text, as speech is inherently continuous and multidimensional, requiring various speech attributes such as acoustic properties, semantic meaning, and contextual clues (Ju et al., 2024). Traditional approaches using feature representations such as Mel-Spectrograms (Sheng et al., 2019), Mel-frequency cepstral coefficients (MFCCs) (Juvela et al., 2018), and Waveforms (Kim et al., 2021) have proven inadequate in capturing this full spectrum of information, resulting in suboptimal performance in downstream tasks such as speech synthesis (Ju et al., 2024).
@@
Figure 1: An overview of speech tokenization approaches using discrete acoustic, semantic, and contextual tokens. DM-Codec integrates these multimodal representations for robust speech tokenization, learning comprehensive speech representations.
@@
While these approaches have shown promising results, they often overlook a crucial aspect of speech representation: the integration of contextual language information. Language models (LMs) have demonstrated a remarkable ability to learn contextual representations that capture the meaning of tokens based on their broader linguistic context (Devlin et al., 2019). These contextual representations can provide essential insights into speech representation, allowing for a more nuanced understanding of words in varying linguistic contexts. Our empirical investigations also reveal that existing discrete speech representation models struggle to align reconstructed speech with accurate textual form, resulting in elevated Word Error Rates (WER) and Word Information Lost (WIL) scores in speech transcription tasks. This observation underscores the need for a more comprehensive approach to speech tokenization that incorporates contextual language information.
@@
To address these challenges, we propose DM-Codec, a novel speech tokenizer that unifies multimodal language and speech representations within a comprehensive tokenizer for speech. Our approach builds on a neural codec architecture incorporating RVQ with encoder, decoder, and discriminator components. Central to our innovation is the introduction of an LM-guided distillation method that effectively incorporates contextual representations into the speech tokenization process. This technique allows DM-Codec capturing the nuances of linguistic context often missed by existing models. Building upon the LM-guided approach, we further propose a hybrid distillation method combining both LM and speech model (SM)-guided techniques. The distillation method only utilizes the LM and SM during training, without increasing model complexity and parameters, and eliminates the need for them during inference. To the best of our knowledge, we are the first to attempt to integrate all three essential aspects of speech representation—acoustic, semantic, and contextual—within a single codec. See Figure 1 for a depiction.
Through extensive experimentation on the LibriSpeech benchmark dataset (Panayotov et al., 2015), we demonstrate the superiority of DM-Codec, which achieves significantly lower WER and WIL compared to state-of-the-art baseline speech tokenizers. Specifically, DM-Codec achieves a WER of $4.05$ and a WIL of 6.61, outperforming SpeechTokenizer ($4.49$, $7.10$), FACodec ($4.68$, $7.33$), and EnCodec ($4.53$, $7.17$). Furthermore, DM-Codec exhibits improved speech quality, as evidenced by its Virtual Speech Quality Objective Listener (ViSQOL) score of $3.26$, surpassing the performance of baseline models (EnCodec: $3.08$; SpeechTokenizer: $3.09$; FACodec: $3.13$).
Our research makes the following key contributions:
@@
  • We introduce DM-Codec, a novel speech tokenizer that incorporates contextual representations via an LM-guided distillation method.
@@
  • We present a novel combined LM and SM-guided representation distillation approach, uniting acoustic, semantic, and contextual representations into a unified framework.
  • Through comprehensive experiments and ablation studies, we demonstrate the effectiveness of DM-Codec in preserving increased contextual information and enhancing the retention of acoustic and speech information in reconstructed speech.
@@ 2 Proposed Method
In this section, we present DM-Codec, a novel speech tokenizer designed to encapsulate a comprehensive fusion of multimodal (acoustic, semantic, and contextual) representations. As illustrated in Figure 2, we propose two distinct training approaches to incorporate these representations: (i) a language model (LM)-guided distillation method, and (ii) a combined LM and self-supervised speech model (SM)-guided distillation method. The first approach distills contextual representations from the LM and integrates them with learned acoustic representations. The second approach combines SM and LM to further incorporate semantic representations with contextual and acoustic representations. It ensures that DM-Codec captures the essential elements of speech by harmonizing the acoustic features with contextual and semantic information. The following subsections detail our proposed distillation methods (§2.1), model details (§2.2), and components (§2.3).
2.1 Speech and Language model Guided Distillation
@@
Our approach first transcribes the raw speech $\mathbf{x}$ into its corresponding text $\mathbf{x^{\prime}}$ using a Speech-to-Text (STT) model $M_{STT}$, such that $\mathbf{x^{\prime}}\,=\,M_{STT}(\mathbf{x})$. For simplicity, we omit any post-processing techniques on the $\mathbf{x^{\prime}}$. Subsequently, we pass the text $\mathbf{x^{\prime}}$ through a pretrained language model $M_{LM}$ to obtain contextual representations of $\mathbf{x^{\prime}}$, tokenized into a set of tokens, $\mathcal{T}=\{t_{i}\}_{i=1}^{n}$. For each token $t_{i}$, we extract its corresponding layer-wise hidden representations $\{\mathbf{h}_{i}^{l}\}_{l=1}^{L}$, where $L$ denotes the total number of layers in $M_{LM}$. We utilize all layer representations to derive the representations for each token, as each layer of a pre-trained language model captures hierarchical and contextually distinct information (Niu et al., 2022; Kovaleva et al., 2019; Hao et al., 2019). To obtain the contextual representation $\mathbf{S}_{i}$ for token $t_{i}$, we average the hidden representations across all layers, yielding $\mathbf{S}_{i}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{h}_{i}^{l}$, where $\mathbf{S}_{i}\in\mathbb{R}^{D}$ where $D$ is hidden dimension. Consequently, we obtain the contextual representations $\mathbf{S}=\{\mathbf{S}_{i}\}_{i=1}^{n}$ for the speech input $\mathbf{x}$, which captures the contextually diverse information from $M_{LM}$.
@@
Simultaneously, we process the raw speech $\mathbf{x}$ through an Encoder $\mathbf{E}(\mathbf{x})$ to obtain the latent feature $\mathbf{v}$. We then pass $\mathbf{v}$ through a Residual Vector Quantizer (RVQ) to obtain quantized features $\mathbf{Q}=\{\mathbf{Q}_{k}\}_{k=1}^{K}$, where $K$ represents the number of quantization layers in the RVQ, and $\mathbf{Q}_{k}\in\mathbb{R}^{D^{\prime}}$ where $D^{\prime}$ is hidden dimension of $k^{th}$ RVQ layer. These quantized features are subsequently used to reconstruct the audio $\mathbf{\hat{x}}$ via a decoder. To align the quantized feature $\mathbf{Q}_{k}$ with the LM distilled features $\mathbf{S}_{i}$, we apply a linear transformation $\mathbf{Q}_{k}^{\prime}=\mathbf{W}\mathbf{Q}_{k}$, where $\mathbf{W}\in\mathbb{R}^{D^{\prime}\times D}$, ensuring the dimensional consistency for the distillation process.
Figure 2: DM-Codec framework consists of an encoder that extracts latent representations from the input speech signal. These latent vectors are subsequently quantized using a Residual Vector Quantizer (RVQ). We designed two distinct distillation approaches: (i) distillation from a language model, and (ii) a combined distillation from both a language model (LM) and a speech model (SM). These approaches integrate acoustic, semantic, and contextual representations into the quantized vectors to improve speech representation for downstream tasks.
@@
LM Guided Distillation: In this approach, we distil the LM representations $\mathbf{S}$. To calculate the LM-guided distillation loss, we adopt a continuous representation distillation technique, similar to the one employed by SpeechTokenizer (Zhang et al., 2024a), which maximizes the cosine similarity at the dimension level across all time steps. In our case, we calculate the continuous representation distillation of the transformed quantized features $\mathbf{Q}_{k}^{\prime}$ and the LM representation features $\mathbf{S}$ as follows:
@@
Here, the notation $(:,d)$ indicates a vector that includes values from all time steps at the $d^{th}$ dimension. The function $\sigma(\cdot)$ represents the sigmoid activation function, commonly used to squash input values into a range between 0 and 1.
@@
Combined LM and SM Guided Distillation: To further enhance the capabilities of DM-Codec, we propose a hybrid approach that utilizes both audio and text modalities. To derive semantic representations from the speech model (SM), we adopt a similar distillation strategy as we used for the LM. We first pass the raw speech $\mathbf{x}$ through the pretrained speech model $M_{SM}$, which generates its own set of layer-wise hidden representations $\{\mathbf{h}_{j}^{l}\}_{l=1}^{L}$. The semantic features are derived by averaging the hidden states across all layers, yielding $\mathbf{A}_{j}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{h}_{j}^{l}$, where $\mathbf{A}_{j}\in\mathbb{R}^{D}$. This process results in the semantic representations $\mathbf{A}=\{\mathbf{A}_{j}\}_{j=1}^{n}$ for the speech input $\mathbf{x}$. The distillation loss in this case considers both the LM and SM representations, jointly optimizing for the alignment of the quantized features $\mathbf{Q}_{k}^{\prime}$ with the representations $\mathbf{A}$ and $\mathbf{S}$ derived from $M_{SM}$ and $M_{LM}$, respectively. Finally, the distillation loss for the SM, $\mathcal{L}_{SM}$, is first computed, followed by averaging with the LM distillation loss, $\mathcal{L}_{L}$, to ensure a balanced contribution from both losses. The combined distillation loss is computed as:
@@
$\mathcal{L}_{SM}$ $=$ $-\frac{1}{D}\sum_{d=1}^{D}\log\left(\sigma\left(\frac{\mathbf{Q}_{k}^{\prime(:,d)}\cdot\mathbf{A}^{(:,d)}}{\|\mathbf{Q}_{k}^{\prime(:,d)}\|\|\mathbf{A}^{(:,d)}\|}\right)\right)$ (2)
$\mathcal{L}_{LS}$ $=$ $\frac{1}{2}\left(\mathcal{L}_{SM}+\mathcal{L}_{L}\right)$ (3)
@@
This formulation ensures that DM-Codec effectively integrates both acoustic and semantic knowledge from SM, along with the contextual information provided by LM, resulting in a more robust and comprehensive set of features for speech discretization.
2.2 Model Details

### v2

@@ Abstract
Recent advancements in speech-language models have yielded significant improvements in speech tokenization and synthesis. However, effectively mapping the complex, multidimensional attributes of speech into discrete tokens remains challenging. This process demands acoustic, semantic, and contextual information for precise speech representations. Existing speech representations generally fall into two categories: acoustic tokens from audio codecs and semantic tokens from speech self-supervised learning models. Although recent efforts have unified acoustic and semantic tokens for improved performance, they overlook the crucial role of contextual representation in comprehensive speech modeling. Our empirical investigations reveal that the absence of contextual representations results in elevated Word Error Rate (WER) and Word Information Lost (WIL) scores in speech transcriptions. To address these limitations, we propose two novel distillation approaches: (1) a language model (LM)-guided distillation method that incorporates contextual information, and (2) a combined LM and self-supervised speech model (SM)-guided distillation technique that effectively distills multimodal representations (acoustic, semantic, and contextual) into a comprehensive speech tokenizer, termed DM-Codec. The DM-Codec architecture adopts a streamlined encoder-decoder framework with a Residual Vector Quantizer (RVQ) and incorporates the LM and SM during the training process. Experiments show DM-Codec significantly outperforms state-of-the-art speech tokenization models, reducing WER by up to 13.46%, WIL by 9.82%, and improving speech quality by 5.84% and intelligibility by 1.85% on the LibriSpeech benchmark dataset. Code, samples, and checkpoints are available at github.com/mubtasimahasan/DM-Codec
DM-Codec: Distilling Multimodal Representations for Speech Tokenization
[…]
@@ 1 Introduction
In recent years, the advent of Large Language Models (LLMs) has revolutionized various domains, offering unprecedented advancements across a wide array of tasks (OpenAI, 2024). A critical component of this success has been the tokenization of input data, enabling vast amounts of information processing (Du et al., 2024a; Rust et al., 2021). Inspired by these breakthroughs, significant attention has shifted towards replicating similar successes in the realm of speech understanding and generation (Défossez et al., 2022; Hsu et al., 2021). However, tokenizing speech into discrete units presents unique challenges compared to text, as speech is inherently continuous and multidimensional, requiring various speech attributes such as acoustic properties, semantic meaning, and contextual clues (Ju et al., 2024). Traditional approaches using feature representations such as Mel-Spectrograms (Sheng et al., 2019), Mel-frequency cepstral coefficients (MFCCs) (Juvela et al., 2018), and Waveforms (Kim et al., 2021) have proven inadequate in capturing this full spectrum of information, resulting in suboptimal performance in downstream tasks such as speech synthesis (Ju et al., 2024).
Figure 1: Overview of speech tokenization using discrete acoustic, semantic, and contextual tokens. DM-Codec integrates these features for robust and comprehensive speech representation.
@@
However, these approaches often overlook a key aspect: the integration of contextual information. Language models (LMs) have demonstrated an ability to learn contextual representations that capture the meaning of tokens from broader context (Devlin et al., 2019). These contextual representations can provide essential insights into speech representation, allowing for a more nuanced understanding of words in varying linguistic contexts. Our empirical investigations also reveal that existing discrete speech representation models struggle to align reconstructed speech with accurate textual form, resulting in elevated Word Error Rates (WER) and Word Information Lost (WIL) scores in speech transcription. This observation underscores the need for a more comprehensive approach to speech tokenization that incorporates contextual information.
@@
To address these challenges, we propose DM-Codec, a novel speech tokenizer that unifies multimodal language and speech representations. Central to our innovation is the introduction of an LM-guided distillation method that incorporates contextual representations into the speech tokenization process. This allows DM-Codec capturing the nuances of linguistic context often missed by existing models. Building upon the LM-guided approach, we propose a distillation method combining both LM and speech model (SM)-guided techniques. Moreover, we introduce a [CLS]-token-based distillation strategy that leverages sequence-level holistic representations from the LM, effectively capturing global contextual information. Our distillation method only utilizes the LM and SM during training, without increasing model complexity and parameters, and are not required during inference. To the best of our knowledge, we are the first to attempt to integrate all three essential aspects of speech representation—acoustic, semantic, and contextual—within a single codec. See Figure 1 for a depiction. In addition, to demonstrate the impact of multimodal representation and generalizability of DM-Codec in downstream tasks, we introduce DM-Codec-TTS, a novel multimodal representation distilled neural codec language model.
Through extensive experiments on LibriSpeech (Panayotov et al., 2015), we show DM-Codec’s superiority. DM-Codec achieves a WER of $4.05$ and WIL of $6.61$, outperforming SpeechTokenizer ($4.49$, $7.10$), FACodec ($4.68$, $7.33$), and EnCodec ($4.53$, $7.17$). It also improves speech quality, with a ViSQOL score of $3.26$ and MOS of $3.72$, surpassing EnCodec ($3.08$, $3.09$), SpeechTokenizer ($3.09$, $3.67$), and FACodec ($3.13$, $3.70$). DM-Codec-TTS excels on LibriSpeech and VCTK, outperforming USLM and Vall-E. On LibriSpeech, it achieves a WER of $5.08$, WIL of $7.32$, MOS of $3.70$, and SMOS of $3.89$, while on VCTK, it achieves a WER of $3.58$, WIL of $5.65$, MOS of $3.78$, and SMOS of $3.85$. Notably, DM-Codec-TTS-small achieves a WER of $10.26$, WIL of $13.79$, MOS of $3.24$, and SMOS of $3.20$, outperforming USLM (libri) across all metrics despite using a smaller dataset.
We summarize our contributions below:
@@
  • We introduce DM-Codec, a novel speech tokenizer that incorporates contextual representations via the LM-guided distillation method.
@@
  • We present a novel combined LM and SM-guided distillation method, unifying acoustic, semantic, and contextual representations.
  • We propose a [CLS]-token-based distillation strategy that captures global contextual information from the LM, facilitating better alignment and transfer of contextual features.
  • We introduce DM-Codec-TTS, a neural codec language model, demonstrating the generalizability of the DM-Codec framework in downstream tasks.
2 Related Work
Tokenization Techniques in Speech. Tokenization in speech processing can be broadly categorized into two main approaches: (i) speech encoder-based and (ii) language-based. In the speech encoder-based tokenization approach, a pretrained speech encoder serves as a teacher model, providing semantically rich audio representations. These representations are then used to guide the training model, either through an alignment network (Messica and Adi, 2024) or by optimizing specific losses (Zhang et al., 2024a; Liu et al., 2024). Language-based tokenization approach involves processing audio through a speech encoder to obtain discrete representations or using the corresponding text to feed into a language model. The representations from the language model are then utilized either to learn a tokenizer for speech or to reconstruct speech (Turetzky and Adi, 2024; Hassid et al., 2024; Zhang et al., 2024b; Wang et al., 2024). Besides, (Zhang et al., 2024b) proposed SpeechLM where two discrete tokenizers were introduced and learned in an unsupervised way and converted the speech and text in a shared discrete space.
Discrete Speech Representation. There are two well-known methods for discrete speech representation: semantic tokens and acoustic tokens. Semantic tokens are derived through self-supervised learning (SSL) techniques for speech (Baevski et al., 2019; Hsu et al., 2021; Chung et al., 2021) and capture abstract, high-level features that relate to general, symbolic aspects of speech, while omitting details related to speaker identity and acoustic characteristics. In contrast, acoustic tokens are obtained using neural audio codecs (Zeghidour et al., 2021; Défossez et al., 2022; Yang et al., 2023) and focus on delivering precise reconstructions of acoustic features. However, recent models (Turetzky and Adi, 2024; Liu et al., 2024; Shi et al., 2024) have shown that speech models based on self-supervised learning (SSL) are effective at extracting acoustic representations where LMs be employed to refine these models further, enhancing their ability to extract more nuanced semantic representations. Recent works (Huang et al., 2024; Du et al., 2024b; Ahasan et al., 2025) focused on improving the quality of discrete speech representation by retaining semantic information for both speech understanding and generation.
Textual Language Models in Speech. Research on speech models, including works by (Nguyen et al., 2023), (Borsos et al., 2023), and (Kharitonov et al., 2022), has focused on utilizing raw audio to extract prosodic features, identify speaker characteristics, and generate audio without depending on textual features or supervision from textual LMs. In contrast, many newer methods have started using audio encoders to transform audio signals into discrete tokens, which can be processed by textual LMs. TWIST method introduced by (Hassid et al., 2024) initializes the weights of the SpeechLM using a pre-trained text LM, showing that this combination significantly improves performance. Similarly, the SELM model developed by (Wang et al., 2024) leverages GPT (Radford, 2018; Radford et al., 2019) as its foundation due to its enhanced parallel processing capabilities and capacity. However, text-based LLMs such as GPT-3 (Brown, 2020) and Llama (Touvron et al., 2023) are essential for speech modeling. Once discrete audio representations are obtained, these large text models are trained to enhance the original text embedding space, as explored in studies by (Zhang et al., 2023), (Fathullah et al., 2023), (Shu et al., 2023), and (Rubenstein et al., 2023). This trend of integrating textual LMs into speech modeling has become increasingly popular in recent research.
Figure 2: DM-Codec framework consists of an encoder that extracts latent speech representations, and quantized using a Residual Vector Quantizer (RVQ). We propose two distillation approaches: (i) from a language model (LM), and (ii) from both an LM and a speech model (SM), integrating acoustic, semantic, and contextual features to enhance speech representations for downstream tasks.
@@ 3 Proposed Method
This section introduces DM-Codec, a novel speech tokenizer that distills multimodal (acoustic, semantic, and contextual) representations. As shown in Figure 2, we propose two training approaches: (i) LM-guided distillation, integrating contextual and acoustic representations, and (ii) combined LM and SM-guided distillation, incorporating semantic, contextual, and acoustic representations. Additionally, a [CLS] token-guided distillation leverages sequence-level holistic contextual representations. We also present DM-Codec-TTS, integrating DM-Codec into a neural codec language model. The following subsections detail the distillation methods (§3.1), DM-Codec model and training objectives (§3.2, §3.3), and DM-Codec-TTS (§3.4).
3.1 Multimodal Representation Distillation
@@
We first transcribe the raw speech $\mathbf{x}$ into its corresponding text $\mathbf{x^{\prime}}$ using an Automatic Speech Recognition (ASR) model $M_{ASR}$, such that $\mathbf{x^{\prime}}\,=\,M_{ASR}(\mathbf{x})$. $M_{ASR}$ serves as an tool for converting speech to text, and provides no supervision or influence on model training. For simplicity, we omit any post-processing techniques on the $\mathbf{x^{\prime}}$. Subsequently, we pass the text $\mathbf{x^{\prime}}$ through a pretrained language model $M_{LM}$ to obtain contextual representations of $\mathbf{x^{\prime}}$, tokenized into a set of tokens, $\mathcal{T}=\{t_{i}\}_{i=1}^{n}$. For each token $t_{i}$, we extract its corresponding layer-wise hidden representations $\{\mathbf{h}_{i}^{l}\}_{l=1}^{L}$, where $L$ denotes the total number of layers in $M_{LM}$. We utilize all layer representations to derive the representations for each token, as each layer of a pre-trained language model captures hierarchical and contextually distinct information (Niu et al., 2022; Kovaleva et al., 2019; Hao et al., 2019). To obtain the contextual representation $\mathbf{S}_{i}$ for token $t_{i}$, we average the hidden representations across all layers, yielding $\mathbf{S}_{i}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{h}_{i}^{l}$, where $\mathbf{S}_{i}\in\mathbb{R}^{D}$ where $D$ is the hidden dimension. Consequently, we obtain the contextual representations $\mathbf{S}=\{\mathbf{S}_{i}\}_{i=1}^{n}$ for the speech input $\mathbf{x}$, which captures the contextually diverse information from $M_{LM}$.
@@
Simultaneously, we process the raw speech $\mathbf{x}$ through an Encoder $\mathbf{E}(\mathbf{x})$ to obtain the latent features $\mathbf{v}$, with sequence length $T^{\prime}$. We then pass $\mathbf{v}$ through a Residual Vector Quantizer (RVQ) to obtain quantized features $\mathbf{Q}=\{\mathbf{Q}_{k}\}_{k=1}^{K}$, where $K$ represents the number of quantization layers in the RVQ, and $\mathbf{Q}_{k}\in\mathbb{R}^{D^{\prime}}$ where $D^{\prime}$ is the hidden dimension of $k^{th}$ RVQ layer. $\mathbf{Q}_{k}$ capture comprehensive speech information across the entire utterance and are subsequently used to reconstruct the audio $\mathbf{\hat{x}}$ via the decoder. The holistic information of the entire speech segment, represented in discrete form within $\mathbf{Q}_{k}$, allows us to directly match it with $\mathbf{S}$ by padding the tokens $n$ to the quantized sequence length $T^{\prime}$. The padded tokens serve only to measure feature dimension similarity rather than impose temporal alignment. To distill contextual information from $\mathbf{S}$ into $\mathbf{Q}_{k}$, we apply a linear transformation $\mathbf{Q}_{k}^{\prime}=\mathbf{W}\mathbf{Q}_{k}$, where $\mathbf{W}\in\mathbb{R}^{D^{\prime}\times D}$, ensuring the dimensional consistency for alignment and updates to occur within each feature dimension without necessitating strict temporal correspondence.
@@
LM Guided Distillation: In this approach, we distil the LM representations $\mathbf{S}$. To calculate the distillation loss, we adopt continuous representation distillation (Zhang et al., 2024a), which maximizes the cosine similarity at the feature dimension axis $D$ with the motivation to focus on similarity within each feature dimension $D$ rather than solely focusing on the overall output similarity. We calculate the continuous representation distillation of the transformed quantized features $\mathbf{Q}_{k}^{\prime}$ and the LM representation features $\mathbf{S}$ as follows:
@@
Here, the notation $(:,d)$ indicates a vector that includes values from all time steps at the $d^{th}$ dimension, where $d\in[1,D]$ in the $D$-dimensional feature space. The function $\sigma(\cdot)$ represents the sigmoid activation function. $\sum_{d=1}^{D}$ computes the loss by aggregating over all feature dimensions, ensuring that updates occur at the feature level rather than at specific time steps.
@@
Combined LM and SM Guided Distillation: We enhance DM-Codec with a hybrid approach using both audio and text modalities. To derive semantic representations from the speech model (SM), we adopt a similar distillation strategy as we used for the LM. We first pass the raw speech $\mathbf{x}$ through the pretrained speech model $M_{SM}$, generating layer-wise hidden representations $\{\mathbf{h}_{j}^{l}\}_{l=1}^{L}$. The semantic features are derived by averaging the hidden states across all layers, yielding $\mathbf{A}_{j}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{h}_{j}^{l}$, where $\mathbf{A}_{j}\in\mathbb{R}^{D}$. This yields semantic representations $\mathbf{A}=\{\mathbf{A}_{j}\}_{j=1}^{n}$ for speech input $\mathbf{x}$. The distillation loss in this case considers both the LM and SM representations, jointly optimizing the quantized features $\mathbf{Q}_{k}^{\prime}$ with the representations $\mathbf{A}$ and $\mathbf{S}$ derived from $M_{SM}$ and $M_{LM}$, respectively. We first calculate the distillation loss for the SM, $\mathcal{L}_{S}$, followed by averaging with the LM distillation loss, $\mathcal{L}_{L}$, to ensure a balanced contribution from both losses, as follows:
@@
$\mathcal{L}_{S}=-\frac{1}{D}\sum_{d=1}^{D}\log\left(\sigma\left(\frac{\mathbf{Q}_{k}^{\prime(:,d)}\cdot\mathbf{A}^{(:,d)}}{\|\mathbf{Q}_{k}^{\prime(:,d)}\|\|\mathbf{A}^{(:,d)}\|}\right)\right)$ (2)
$\mathcal{L}_{LS}=\frac{1}{2}\left(\mathcal{L}_{L}+\mathcal{L}_{S}\right)$ (3)
@@
This formulation ensures that DM-Codec integrates acoustic representation learned by the codec architecture with semantic knowledge from SM and contextual knowledge from LM.
[CLS] Token Guided Distillation: We introduce a [CLS] token-guided distillation strategy, leveraging the [CLS] token’s sequence-level holistic representation to capture global contextual information from LM. This approach eliminates the need for fine-grained temporal alignment while preserving essential linguistic features. Moreover, CLS-guided distillation employs sentence-level knowledge transfer, which has been shown to be more robust in noisy settings (Wei et al., 2024). For this method, we use the layer-wise hidden representations of the [CLS] token alone. These representations, averaged across all layers, are denoted as $\mathbf{S}_{\text{[CLS]}}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{h}_{\text{[CLS]}}^{l}$, where $\mathbf{S}_{\text{[CLS]}}\in\mathbb{R}^{D}$. To match the sequence length $T^{\prime}$ of the quantized features $\mathbf{Q}_{k}^{\prime}$, the [CLS] token representation is repeated $T^{\prime}$ times, yielding $\mathbf{S}^{\prime}=\{\mathbf{S}_{\text{[CLS]}},\mathbf{S}_{\text{[CLS]}},\ldots,\mathbf{S}_{\text{[CLS]}}\}$.
The distillation loss follows the same formulation as the LM-guided distillation loss (Eqn. 1), replacing $\mathbf{S}$ with $\mathbf{S}^{\prime}$. This enables the distillation process to leverage the global contextual information encoded in the [CLS] token while ensuring greater alignment with the sequence length $T^{\prime}$.
3.2 DM-Codec: Model Details